OpenStudy (anonymous):

my answer is b Solve log4 x - 1 = 2 (1 points) Answer x = 64 x = 17 x = 4 x = 2

OpenStudy (johnweldon1993):

correct :)

sam (.sam.):

This? \[\log _4(x)-1=2\]

sam (.sam.):

or \[\log (4x)-1=2\]

sam (.sam.):

I guess is the first one, b is not correct

OpenStudy (johnweldon1993):

@.Sam. she means log base 4.....it corresponds with her other questions

OpenStudy (anonymous):

the first one i think, it's not very clear

OpenStudy (anonymous):

is B right?

sam (.sam.):

nope, try again :)

OpenStudy (johnweldon1993):

\[\log_{4}(x-1) = 2 \] I believe she has it like this

sam (.sam.):

i see

sam (.sam.):

correct then :)

OpenStudy (anonymous):

yayy!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@Sam can i try? hihi

OpenStudy (anonymous):

give me the answer, please. I want it, too

OpenStudy (johnweldon1993):

haha good job @InsanelyChaotic :)

OpenStudy (anonymous):

thank you!!! @johnweldon1993